# MIGE: A Unified Framework for Multimodal Instruction-Based Image Generation and Editing

### v1

@@ 4.3.2 Effectiveness of Feature Fusing
MIGE employs a feature fusion mechanism in the multimodal encoder to integrate semantic features from ViT and visual features from VAE. As shown in Table 1 and Table 2, compared to the model without VAE features (denoted as “wo_VAE feature”), incorporating VAE features significantly improves detail preservation in reference images, benefiting both subject-driven image generation and instruction-based image editing. This is particularly evident in the improved CLIP-I and DINO scores and the significant reduction in L1 and L2 metrics, demonstrating that the inclusion of additional visual features helps maintain consistency between the input and output.
@@
4.3.3 Effectiveness of Instruction-based Subject-driven Image Editing Dataset
@@
Joint training in subject-driven image generation and instruction-based image editing enables generalization to instruction-based subject-driven image editing (denoted as “subject data + edit data”). To enhance the capability of MIGE in this new task, particularly in understanding spatial terms and size descriptions, we constructed a task-specific dataset for joint training. As shown in Table 3(a) and Table 3(b), the task-specific data significantly improved the model’s overall abilities. This demonstrates the effectiveness of our constructed dataset, and the proposed data generation pipeline serves as a valuable reference for future dataset construction.
4.3.4 Effectiveness of Multimodal Instruction
@@
Existing instruction-based editing works Brooks et al. (2023); Sheynin et al. (2024); Zhang et al. (2024) typically use text instructions as the conditional input, while we extend this to multimodal instructions. To measure the benefit of multimodal instructions, we trained the model with text-only editing instructions for comparison. As shown in Table 2, using multimodal instructions consistently improves performance over text-only instructions (denoted as “wo_multimodal instruction”). This enhances input-output consistency and instruction-following ability in multi-task training. The significant improvement in the L1 and L2 metrics indicates finer control over images and more accurate edits. While text-only instructions provide the necessary changes, the high CLIPdir score and lower values in other metrics show that multimodal instructions add visual context, enabling more precise and faithful modifications.
5 Conclusion
@@
We present MIGE, a unified framework that combines subject-driven generation and instruction-based editing. By leveraging multimodal instructions and conditional input, MIGE enables joint training that reinforces both tasks. This approach improves task synergy and addresses data scarcity. Joint training also unlocks new capabilities like instruction-based subject-driven image editing. We introduce pipelines for this new task to construct training data and MIGEBench for evaluation. Our experiments show that joint training leads to significant improvements in subject fidelity and instruction adherence, demonstrating the effectiveness of unifying these tasks. This integration enhances controllability and offers promising directions for future multimodal image generation and editing.
6 Limitations
Our model faces common challenges in context accuracy, complex composition, and visual faithfulness, especially in multi-entity tasks. Ethical concerns, like deepfakes, also arise. Nevertheless, our framework shows promise in enabling unified controlled image generation. Instruction-based subject-driven generation remains challenging, with difficulties in handling spatial relationships and adjusting subject sizes to fit the background. Additionally, the technology carries the risk of generating misleading images, but our work is intended solely for research, emphasizing its exploratory nature rather than real-world deployment.
Appendix A Benchmark Construction
Instruction-based subject-driven image editing is a novel task lacking evaluation benchmarks. Previous works Li et al. (2023b); Yang et al. (2024) did not support instructions and lacked ground truth target images. To address this, we designed a benchmark construction pipeline with multiple manual inspections to ensure image and caption quality.
As shown in Figure 7, we first filter valid editing pairs using a Gradio Abid et al. (2019) interface, as shown in Figure 8. Then, Qwen2.5-14B-Instruct extracts entity names and constructs multimodal instructions. After manually reviewing object names, we use Grounding DINO Liu et al. (2025) for bounding boxes and SAM for segmentation. Cropped entities are saved with black and white backgrounds for diverse test scenarios. Prompts are detailed in Appendix E.
For evaluation, Qwen2-VL-7B generates captions that compare the generated images with their ground truth counterparts. These captions are then manually reviewed to identify and remove any unrecognized entities. This review process ensures that the generated results are accurate and reliable. The final benchmark consists of 500 samples for both subject addition and replacement tasks. By carefully avoiding artifacts from inpainting and incorporating multiple rounds of manual review, we guarantee the high quality of the data. Figures 9 and 10 display some example pairs from the benchmark to illustrate the evaluation process.
Figure 7: The pipeline of benchmark construction.
Figure 8: The annotation interface for benchmarks is built with Gradio.
Figure 9: Subject addition examples in our benchmark.
Figure 10: Subject replacement examples in our benchmark.
Appendix B Dataset Construction
To enable joint training of MIGE, we designed a series of pipelines for data construction and processing. Our dataset includes three tasks: subject-driven image generation, instruction-based image editing, and instruction-based subject-driven image editing. The proportions of the training data are shown in Figure 11. In this section, we will detail the pipelines for processing each data component and other related details.
Figure 11: Composition of training data.
Subject-driven Image Generation Data Construction
We construct multimodal instruction-image pairs by replacing entities in image captions with corresponding images. The dataset is sourced from BLIP3-GROUNDING-50M Xue et al. (2024) and internal data. For the former, we discard low-quality pairs with CLIP-T scores below 0.255. Entities are filtered based on annotated bounding boxes, retaining those with sizes between 0.05 and 0.8. After segmenting entities using SAM Kirillov et al. (2023), we further compute the CLIP-T score between “a photo of [entity]” and the segmented image to mitigate incomplete pairs. For internal data, we follow the KOSMOS-G and UNIMO-G processing pipelines, using MPT-7B-Instruct Team et al. (2023) to extract entities. The original image and entity text are input into Grounded SAM to obtain entity images, which are filtered by a CLIP-T score above 0.255 and a size ratio between 0.1 and 0.85. To enhance subject preservation, we incorporate 112,846 samples from subject200k, as in OminiControl Tan et al. (2024).
Instruction-based Image Editing Data Construction
Previous works on instruction-based editing relied solely on text, covering object-level additions, deletions, modifications, and global changes in background or style. We integrated data from InstructPix2Pix Brooks et al. (2023), UltraEdit Zhao et al. (2024), MagicBrush Zhang et al. (2024), SEED-Data-Edit Ge et al. (2024), GIER Shi et al. (2020), and HQ-Edit Hui et al. (2024), filtering out instances with target image aesthetic scores above 5.5. To enhance multimodal instruction alignment, we reformulated text-based instructions based on task type. Specifically, for instructions containing phrases like “in the image” or “to the image,” we replaced them with “in<imagehere>” or “to<imagehere>.” For others, we appended “of<imagehere>” for global edits or “in<imagehere>” for localized modifications.
Instruction-based Subject-driven Image Editing Data Construction
To enhance instruction-based subject-driven image editing, we propose dedicated data construction pipelines. For subject addition, we design a pipeline inspired by SAM-FB He et al. (2024) to construct background images, foreground entities, and multimodal instructions, as shown in Figure 4(a). We start with filtering SA-1B Kirillov et al. (2023) images with an aesthetic score above 5 and select entities where the foreground-to-image ratio is between 0.1 and 0.5. Qwen2-VL-7B Wang et al. (2024) is used to verify entity completeness. The selected entities serve as foregrounds, while the backgrounds are inpainted using LAMA Suvorov et al. (2022). We then extract foreground-related text using Qwen2-VL-7B and generate multimodal instructions with GPT-4o Hurst et al. (2024). Due to resource constraints, we process the first 500,000 SA-1B samples, yielding 193,247 multimodal instruction-foreground-background pairs.
For subject replacement, as shown in Figure 4(b), we aim at constructing foreground entities and multimodal instructions. We filter SEED-Data-Edit’s part1-unsplash based on predefined rules, retaining only replace-task data where the match score in the annotations exceeds 0.3 and the target image’s aesthetic score is above 5.5. Qwen2.5-14B-Instruct Team (2024) identifies whether the main entity is being replaced. The target entity is replaced with <imagehere> to construct a multimodal prompt. Using the annotated edited subject information, we apply Grounded SAM for segmentation, obtaining 79,693 multimodal instruction-replace entity image-source image-target image pairs. In addition, we construct 34,947 virtual try-on samples using IDM-VTON Choi et al. (2024).
Appendix C Evaluation Details
We evaluated the subject-driven generation capability of MIGE on DreamBench, which contains 750 prompts covering 30 subjects. For each prompt, four images were generated using seeds 0, 1, 2, and 3, resulting in a total of 3,000 images. Following KOSMOS-G, we selected one image per prompt. Then we extracted the subject using Grounded SAM and used it as input, aligning with the training process. Subject fidelity was assessed using DINO and CLIP-I, while CLIP-T measured adherence to multimodal instructions.
For instruction-based editing evaluation, the Emu Edit benchmark contains known issues, including incorrect image-caption pairs and duplicate source-target captions. Prior works have handled these inconsistencies differently, leading to incomparable results. Therefore, we reimplemented all currently available open-source methods on both test sets using a fixed random seed (seed=0) for consistency. Consequently, CLIP-I, L1, and DINO were computed on 3,589 Emu Edit test samples, while CLIPdir and CLIP-T were evaluated on 2,899 samples after removing problematic data.
Appendix D Related Work
Subject-Driven Image Generation
Recent advances in generative models have significantly improved image customization, enabling finer and more flexible control over generated content. Subject-driven image generation, which preserves the features of given images while creating new content, has garnered increasing attention. DreamBooth Ruiz et al. (2023) fine-tunes the entire U-Net using multiple images of the same subject, making it resource-intensive. IP-Adapter Ye et al. (2023) introduces a test-time tuning-free approach by decoupling image and text, training only additional attention layers, but struggles to balance prompt adherence and image fidelity. FastComposer Xiao et al. (2024a) and Blip-Diffusion Li et al. (2024a) leverage embeddings to condition images on text but face challenges in subject preservation and efficiency with generalized vision-language inputs. KOSMOS-G Pan et al. (2023) and UNIMO-G Li et al. (2024b) use Multimodal Large Language Models (MLLMs) as multimodal encoders for greater input flexibility but require large datasets to train task-specific encoders and alignment networks. However, these methods lack strong instruction-following capabilities, limiting their scalability to new tasks and applications.
Instruction-based Image Editing
Instruction-based image editing modifies images at the object level (e.g., adding, removing, or altering objects) or the image level (e.g., changing style or environment) based on user instructions. This approach offers greater flexibility than caption-based methods but demands stronger instruction-following capabilities. Inst-inpaint Yildirim et al. (2023) focuses on object removal, while EraseDraw Canberk et al. (2025) fine-tunes a model for object addition using data derived from removal tasks. InstructPix2Pix Brooks et al. (2023), trained on a large synthetic dataset, supports four types of edits, and MagicBrush Zhang et al. (2024) refines this capability by fine-tuning on real, manually annotated data. EmuEdit Sheynin et al. (2024) enhances diffusion-based editing with a task vector to differentiate tasks. However, these methods often struggle with maintaining input-output consistency.
Figure 12: Qualitative results of subject-driven image generation (top) , instruction-based image editing (middle), and instruction-based subject-driven image editing (bottom).
Appendix E Prompts
LLMs and MLLMs are integral to our training data and benchmark construction pipeline, and the prompts used are presented in this section.
Prompt for Qwen2-VL-7B to determine the completeness of subject in the image
Prompt: Determine if the subject in the image is complete. If it is complete and not an abstract object such as background, grass, sky, tree, stone, or part of another item, please return True. Otherwise, return False.
Table 4: Prompt design for subject completeness evaluation in an image. The model should identify whether the subject is complete, not abstract elements.
Prompt for GPT-4o to generate multimodal instruction
Prompt: The object is {object_name}. It is located in a bounding box with coordinates ({x}, {y}, {w}, {h}) on an image of size {width}x{height}. Describe its size, relative position, and relation to surrounding objects. Avoid describing the overall scene or unrelated elements.
Your response should start with “Add <imagehere> to the [position] of <imagehere>”. (The first <imagehere> indicates the object and the second indicates the image.) Keep the <imagehere> symbol in the first sentence in your reply.
Answer briefly in two sentences:
Table 5: Prompt design for GPT-4o to generate multimodal instruction for subject addition training data.
Prompt for Qwen2.5-14B-Instruct to determine whether an edit instruction pertains to the main subject of an image
Prompt: You are an assistant that determines whether an edit instruction pertains to the main subject of an image.
The main subject refers to humans or animals only. If the edit instruction is related to the main subject, respond with ’yes’.
If it pertains to background, large areas of vegetation, or environmental information, respond with ’no’.
Here are some examples:
1. replace the grass with sand $\rightarrow$ no
2. replace the trees with palm trees $\rightarrow$ no
3. replace the dirt road with a cobblestone path $\rightarrow$ no
4. replace the bird with a parrot $\rightarrow$ yes
Now, analyze the following instruction and respond accordingly: {instruction}
Table 6: Prompt design for determining if an edit instruction pertains to the main subject of the image. The model evaluates whether the instruction relates to humans, animals, or other environmental features.
Prompt for extracting the main subject and modifying the edit instruction
Prompt 1: Replace Examples
You are an assistant that determines the main subject of an edit instruction and outputs the extracted main subject along with a modified prompt.
Given an edit instruction like “replace X with Y”:
- Extract “Y” (the replacement subject).
- Output the extracted subject.
- Modify the instruction by replacing “Y” with <imagehere> in the original prompt and output the new instruction.
Examples:
1. replace one woman with a man $\rightarrow$ Output: a man, replace one woman with <imagehere>
2. replace the castle with another castle $\rightarrow$ Output: castle, replace the castle with <imagehere>
3. replace the desk with a white one in the bottom middle $\rightarrow$ Output: desk, replace the desk with <imagehere> in the bottom middle
Prompt 2: Add Examples
You are an assistant that extracts the added subject in the edit instruction.
Given an edit instruction like “add X to the image”:
- Extract “X” (the added subject).
- Output the extracted subject.
- Modify the instruction by replacing “X” with <imagehere> in the original prompt and output the new instruction.
Here are some examples:
1. add human over the stone in the bottom left $\rightarrow$ Output: human, add <imagehere> over the stone in the bottom left
2. add a car on the road at the bottom $\rightarrow$ Output: a car, add <imagehere> on the road at the bottom
3. Add an owl on the left shoulder $\rightarrow$ Output: an owl, Add <imagehere> on the left shoulder
Table 7: Prompt designs for extracting and modifying edit instructions based on either replacement or addition of subjects.

### v3

@@ 4.3.2. Effectiveness of Feature Fusing
MIGE employs a feature fusion mechanism in the multimodal encoder to integrate semantic features from ViT and visual features from VAE. As shown in Table 2, Table 1 and Table 3, compared to the model without VAE features (denoted as “wo_VAE feature”), incorporating VAE features significantly improves detail preservation in reference images, benefiting both subject-driven image generation and instruction-based image editing. This is particularly evident in the improved CLIP-I and DINO scores and the significant reduction in L1 and L2 metrics, demonstrating that the inclusion of additional visual features helps maintain consistency between the input and output.
@@
4.3.3. Effectiveness of Instruction-based Subject-driven Image Editing Dataset
@@
Joint training of subject-driven image generation and instruction-based image editing enables generalization to instruction-based subject-driven image editing (denoted as “subject data + edit data”). To enhance the capability of MIGE in this new task, particularly in understanding spatial terms and size descriptions, we constructed a task-specific dataset for joint training. As shown in Table 3(a) and Table 3(b), the task-specific data significantly improved the model’s overall abilities. This demonstrates the effectiveness of our constructed dataset, and the proposed data generation pipeline serves as a valuable reference for future dataset construction.
4.3.4. Effectiveness of Multimodal Instruction
@@
Existing instruction-based editing works (Brooks et al., 2023; Sheynin et al., 2024; Zhang et al., 2024) typically use text instructions as the conditional input, while we extend this to multimodal instructions. To measure the benefit of multimodal instructions, we trained the model with text-only editing instructions for comparison. As a unified model, MIGE demonstrates exceptional performance across various image generation and editing tasks, even outperforming existing task-specific models. As shown in Table 1, using multimodal instructions consistently improves performance over text-only instructions (denoted as “wo_multimodal instruction”). This enhances input-output consistency and instruction-following ability in multi-task training. The significant improvement in the L1 and L2 metrics indicates finer control over images and more accurate edits. While text-only instructions provide the necessary changes, the high CLIPdir score and lower values in other metrics show that multimodal instructions add visual context, enabling more precise and faithful modifications. For more qualitative results, see Supplementary Material.
4.3.5. Comparison of Data Volume
Task-specific models treat tasks separately and typically require large amounts of high-quality data. In contrast, by exploiting the relationships between tasks and the unified structure, MIGE achieves excellent results with minimal data through joint training. As shown in Table 4, MIGE demonstrates that both the model structure and joint training are key to achieving this performance with minimal data, using fewer data than task-specific models and other unified models.
Table 4. Comparison of Data Volume Across Methods
      Methods       Data Volume
             Task-specific       Models       MADD (He et al., 2024a)       3M
      PBE (Yang et al., 2023)       16M
      KOSMOS-G (Pan et al., 2023)       3159.3M
      UNIMO-G (Li et al., 2024c)       304.3M
             Unified       Models       PixWizard (Lin et al., 2024)       30M
      UniReal (Chen et al., 2024c)       100M
      ACE (Han et al., 2024)       700M
      Omnigen (Xiao et al., 2024)       326.06M
      MIGE (ours)       2.28M
5. Conclusion
@@
We present MIGE, a unified framework that combines subject-driven generation and instruction-based editing, leveraging multimodal instructions and conditional input for joint training that enhances task synergy. Joint training also unlocks new capabilities like instruction-based subject-driven image editing. We introduce pipelines for this new task to construct training data and MIGEBench for evaluation. Our experiments show that joint training leads to significant enhancement in subject fidelity and instruction adherence, demonstrating the effectiveness of unifying these tasks. This integration enhances controllability and offers promising directions for future multimodal image generation and editing.